OpenStudy (anonymous):

Find the roots of the quadratic equation x2 − 8x + 8 = 0 . x1 = x2 =

OpenStudy (radar):

\[(-b \pm \sqrt{b ^{2}-4ac)}/2a\]

OpenStudy (radar):

b=-8, a=1, and c=8

OpenStudy (anonymous):

I prefer: \[x = \frac{-b \pm \sqrt{b^2-4ac}}{2a} \] \.[x = \frac{-b \pm \sqrt{b^2-4ac}}{2a} \]

OpenStudy (radar):

that's more betterr

OpenStudy (anonymous):

6.828, 1.17157

OpenStudy (radar):

Car2021, You have everything you need to solve it, including some answers that you can use as a check.

OpenStudy (anonymous):

ok so if i have \[8\pm 4\sqrt{2}/2\]..how would i solve that?

OpenStudy (radar):

You could simplify further and get \[4\pm 2\sqrt{2}\]

OpenStudy (radar):

Solve for both roots

OpenStudy (radar):

Use the + value for one root, then use the - for the other \[4-2\sqrt{2}\]

OpenStudy (radar):

\[4+2\sqrt{2}\]

OpenStudy (anonymous):

soo would it be 2 sqrt 2 for one and 6 sqrt 2 for the other

OpenStudy (radar):

Almost, first solve for the 2 times the square root of 2. this should give you a value of 2.828. Now with that you will use to get the two roots. First add that to 4, that will be one root. Then use it again subtracting it from four getting the second root.

OpenStudy (radar):

Let me know what you get.

OpenStudy (anonymous):

if, ax^2 + bx + c is a quadratic equation then we can find its roots by using a formula http://upload.wikimedia.org/math/3/e/a/3ea647783b5121989cd87ca3bb558916.png now try to solve it yourself. don't just go for the solution.

OpenStudy (radar):

He has solved it down to the point of\[4\pm 2\sqrt{2}\]

OpenStudy (anonymous):

radar ..when u get done will you help me?

OpenStudy (anonymous):

or chatindia..lol

OpenStudy (radar):

Yes, would be glad too, providing I can. I am just reviewing math from the past lol

OpenStudy (anonymous):

lol okay thank you! :))

OpenStudy (radar):

Is your problem already posted laurenlynn?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it's the one that says PLEASE HELP ME!!! at the top..haha

OpenStudy (radar):

O.K. I will look for it.

OpenStudy (radar):

Car2021, I believe you have this problem whipped, cantorset has given you the actual answers in an early post of this thread. Good luck

OpenStudy (anonymous):

what's the problem you are getting in solving this question. this latex code doesn't display in my phone. i can't understand this latex codes.